# Variational Signal Separation for Automotive Radar Interference Mitigation

### v1

@@ Abstract
Algorithms for joint mutual interference mitigation and object parameter estimation are a key enabler for automotive applications of frequency-modulated continuous wave (FMCW) radar. The underlying signal model poses a challenge for signal separation, since both the coherent radar echo and the non-coherent interference influenced by individual multipath propagation channels must be considered. In particular, under certain assumptions, the model is described as a superposition of multipath channels weighted by parametric chirp envelopes in the case of interference.
In this paper, we introduce a method inspired by sparse Bayesian learning (SBL) to detect and estimate radar object parameters while also estimating and successively canceling the interference signal. An augmented probabilistic model is employed that uses hierarchical Gamma-Gaussian prior model for each multipath channel separately. Based on this model an iterative inference algorithm is derived using the variational expectation-maximization (EM) methodology. The algorithm is statistically evaluated in terms of object parameter estimation accuracy and robustness, indicating that it is fundamentally capable of achieving the Cramer-Rao lower bound (CRLB) with respect to the accuracy of object estimates and it closely follows the radar performance achieved when no interference is present.
@@ I-B Contributions and Notations
Although the existing literature is extensive, we identify the following open topics for which our work provides novel insights. Previously proposed algorithms are applied as preprocessing in a conventional range-Doppler processing chain. This approach is practical, but has inherent limitations. In general, the goal of radar sensing is to provide estimates of object parameters, which is equivalent to estimating the parameters of a multipath propagation channel containing static (such as buildings, traffics signs or trees) and moving objects (such as vehicles, bicycle or pedestrians). Note that parametric channel models typically represent multipath propagation as a linear superposition of weighted Dirac delta distributions - or spectral lines - with distinct supports in the underlying dispersion domain (range, angle of arrival, angle of departure, Doppler frequency, and combinations thereof). Therefore, estimating multipath parameters can be cast as a line spectral estimation (LSE) problem [56].
To achieve robust inference performance, it is necessary to consider the entire problem of parameter estimation under the influence of interfering signals. Previous methods are also often based on heuristics or require the prior setting of certain algorithm parameters. This makes it necessary to manually fine-tune the range-Doppler processing chain and makes it difficult giving performance guarantees. Our proposed algorithm is derived within a rigorous theoretical framework. Specific contributions are the following:
@@
  • We propose a novel inference algorithm inspired by SBL [57, 58, 59] for LSE that is able to infer the superposition of sparse line spectra. The proposed algorithm, which is based on the variational EM approach, provides super-resolution by jointly estimating the objects’ range and Doppler parameters, as well as the parameters of the interference multipath channel and signal parameters. This makes it explicitly robust to mutual interference.
@@
  • We include investigations on difficult-to-handle scenarios of model mismatch and poor signal separability. We point out the relative resilience of the algorithm to these challenges, and provide an initial proof-of-concept based on measurement data.
@@ II-A FMCW Radar Under Interference
where we defined the parametrized rectangular function template $\bar{x}(t;T)$ and the chirp $u(t;f_{0},k)$. $A$ is some complex transmit amplitude, $P$ is the number of chirps, $f_{0}$ is the chirp start frequency, $k$ is the chirp slope and $T_{\mathrm{p}}$ is the pulse duration which consists of the actual transmit chirp duration $T_{\mathrm{sw}}$ plus an additional idle time. [^2] Note that actual transmit signals are real-valued but a complex receiver architecture is assumed here. Thus, we immediately use a complex representation for mathematical simplicity. However, it is possible to model a real receiver architecture as well.
@@ III-D Algorithm Implementation
We have now derived structured mean-field variational EM update steps for inference of our model. As mentioned, these steps are interdependent and therefore to be applied iteratively. Explicit forms of the update equations are listed in Appendix B. The algorithm converges to an optimum of the ELBO corresponding to an approximate solution of our signal separation problem. However, it does not necessarily converge to the global optimum, nor is any rate of convergence guaranteed. Our signal separation problem is inherently ambiguous, with many possible combinations of estimates reconstructing the received signal. In addition to the specific update schemes presented previously, the overall scheduling of steps as well as the initialization significantly influence the behavior of the algorithm.
@@
Received signal $\bm{r}$; initial estimates $\hat{\bm{\Phi}}$, $\hat{\bm{\alpha}}$, $\hat{\bm{C}}_{\alpha}$, $\hat{\bm{\gamma}}_{\alpha}$, $\hat{\bm{\Psi}}$, $\hat{\bm{\beta}}$, $\hat{\bm{C}}_{\beta}$, $\hat{\bm{\gamma}}_{\beta}$, $\hat{\lambda}$; thresholds $T_{\alpha}$, $T_{\beta}$
@@
For notational simplicity we have consistently omitted iteration counters from Algorithm 1. Within the sequence of update steps, it uses the most recent values of other estimates. The main iteration consists of two separated subroutines for the estimations of the respective object and interference signals, similar in structure and connected by computations of $\hat{\bm{r}}_{\alpha}$, $\hat{\bm{r}}_{\beta}$, as well as $\hat{\lambda}$. I.e., the estimated residuals are used at every iteration as inputs to the respective interference and object signal estimation subroutines. This scheme is obtained as the interference and object signals are additive and their components factorize according to (14) and (20). Furthermore, as interference is modeled independently over the ramps, separate steps over the index $p$ are applied.
@@
The estimation subroutines of Algorithms 2 and 3 are based on the repeated addition and then update of components by applying (22) and Algorithm 4. The routine adds components it can “find” given the selected threshold, which are then refined or pruned using the now updated estimate of the weight covariance matrix. Hence, the proposed algorithm structure envisions a bottom-up initialization scheme, i.e., that the dictionaries $\hat{\bm{\Phi}}$ and analogously $\hat{\bm{\Psi}}$ are initialized as empty, which is computationally advantageous. The noise precision is initialized as $\hat{\lambda}=(2(NP-1))/{\|\bm{r}\|}^{2}$, i.e., half the received signal power is assumed to be from noise initially.
@@ IV-A Fundamental Object Estimation Performance
Parameter Simulation I Simulation II Measurement I Measurement II
@@ IV-C Statistical Performance Comparison
Figure 5: Statistical results of multi-object simulation scenario. (a)-(c) Evaluated over varying SNR at a fixed SIR of $-20\mathrm{dB}$. (d)-(f) Evaluated over varying SIR at a fixed SNR of $60\mathrm{dB}$.
@@
    mca - using the concept and model of [49]. Our implementation follows that of [72, Algorithm 30].
Fig. 5 illustrates the mean of the resulting metrics computed from 500 realizations. Over SNR, interference leads to a lower limit on the achievable estimation error and the number of misdetections and causes false alarms. With preprocessing, the estimation error and number of misdetections improve, but the performance is still limited and improvements come at the cost of false alarms. This behavior can be explained by considering the effects of preprocessing on the form of the resulting signal. Zeroing merely exploits the time-limited nature of the interference, and hence leads to a signal with missing clusters of samples. MCA considers a signal separation problem, but does not employ a parametric model for the interference bursts and only applies single-ramp on-grid LSE for the object signal. Preprocessing therefore typically causes a distortion of the object signal that is a model error for the subsequent detector, primarily leading to false alarms. Our proposed method mitigates this behavior and reaches performance very close to the No Interference case. The estimation error and the number of misdetections decrease with increasing SNR, and the number of false alarms is constant according to the set detection threshold.
The bottom row of Fig. 5 shows the results at a high SNR, over SIR levels ranging from highly interfered to weakly interfered object signals. Ideally, perfect interference cancellation leads to values identical to the No Interference case at $\mathrm{SNR}=60\mathrm{dB}$. The proposed method approximately achieves this up until $\mathrm{SIR}\approx 0\mathrm{dB}$. Its performance then slightly decreases with decreasing interference power up until $\mathrm{SIR}\approx 15\mathrm{dB}$, after which it slowly recovers. This is the same effect as discussed in Section IV-A. Note that unlike for the simulation scenario of Fig. 3, in the scenario of Fig. 5 the condition of (27) is met roughly at $5\mathrm{dB}<\mathrm{SIR}<10\mathrm{dB}$. This strengthens our hypothesis on the significance of this condition. For the other evaluated methods, we see a bounded improvement in RMSE and misdetection rate at the cost of further false alarms. Naturally, perfect zeroing is independent of interference power while MCA performance is generally improved with increasing SIR.
@@
IV-D Measurement Example
Figure 7: Example of algorithm applied to real automotive radar measurement data. (a) Photo of the measurement scene. (b)-(c) Range-Doppler spectrum of the measured signal without/with interference and corresponding object detections. (d) Range-Doppler spectrum of the interference-cancelled signal and corresponding object detections as a result of the proposed algorithm.
In order to illustrate the practical application of the proposed method, we apply it to radar measurement data from a real traffic scenario. Details on the measurement campaign can be found in previous works [73, 74]. Measurement parameters are summarized in Table I (Measurement I). Note that there is no ground truth of the channel parameters available, as the measurements were done with a single radar sensor in completely uncontrolled scenarios. We select a single measurement and showcase the results qualitatively. Furthermore, the measurements do not inherently contain any interference. In prior work, simulated interference was added to the measured signals. In this work however, measured interference is added, as described in Appendix C, with parameters listed in Table I (Measurement II). The interference was obtained from a separate measurement, due to the difficulty of generating and measuring interference separately in the same inner city traffic environment. Nevertheless, the applied scheme enables a qualitative comparison of our mitigation result to an interference-free one, with both object and interference components obtained using actual radar sensors.
Fig. 7a shows a photo of the selected Measurement I scene. Particularly, visible on the photo is the car in front and to the left of the ego-radar, as well as the facade of a row of buildings to the right. The interference-free spectrum in Fig. 7b hence contains a few close object components with a positive relative velocity, i.e., the car moving away from the sensor. The buildings lead to a large number of components distributed over distances, with negative relative velocities with respect to the measurement vehicle. Other detected components are reflections not easily associated with the camera image or false alarms due to noise. When strong interference is introduced, the range-Doppler spectrum becomes dominated by it, see Fig. 7c. Directly applying object detection onto this signal leads to no discernibly useful result. Finally, the spectrum of the measurement when subtracting the interference estimated by the proposed processing algorithm, and the estimated object components are shown in Fig. 7d. The resulting spectrum is visually very close to the interference-free one and the detected components correspond well to the significant parts of the described scene. This very promising result is nevertheless only preliminary, and a larger scale analysis with better known hardware and measurement environment is a necessary part of future work.
We also qualitatively illustrate the fast-time signal of a single ramp in Fig. 8 as an example. The top graph shows a large interference burst not too dissimilar to, but distinct from the simulated ones of Figs. 2 and 4. The bottom plot provides a close-up view of the signal after subtraction of the interference estimate as well as shows the object signal estimate.
@@ V Conclusion
Specific choices in the proposed algorithm design, initialization and scheduling are essential for its performance. This includes employing grid-less estimation for the object channel, yielding super-resolution accuracy, while using an on-grid algorithm for the interference channels. Parameter optimization and inference of the respective channel weights are derived from a joint update, which decreases the probability of converging to a local optimum of the inherently highly ambiguous signal separation problem.
@@ Appendix B Mean-field Variational EM Updates
We proceed to summarize the solution steps needed to obtain the update equations for our model. Writing out (14), the joint PDF reads
$\log p$ ${(\bm{r},\bm{\alpha},\bm{\gamma}_{\alpha},\bm{\beta},\bm{\gamma}_{\beta},\lambda;\bm{\zeta},\bm{\theta})}$
@@
$\,\quad-\lambda{(\bm{r}-\bm{\Phi\alpha}-\bm{U\Psi\beta})}^{\mathrm{H}}(\bm{r}-\bm{\Phi\alpha}-\bm{U\Psi\beta})$
$\,\quad+(e_{0}-1)\log\lambda-f_{0}\lambda+e_{0}\log f_{0}-\log(\Gamma(e_{0}))$
$\,\quad+\sum_{l}\Bigl{(}-\log\pi+\log\gamma_{\alpha,l}-{|\alpha_{l}|}^{2}\gamma_{\alpha,l}$
$\,\quad\quad+(a_{0}-1)\log\gamma_{\alpha,l}-b_{0}\gamma_{\alpha,l}$
$\,\quad\quad+a_{0}\log b_{0}-\log(\Gamma(a_{0}))\Bigr{)}$
$\,\quad+\sum_{p}\sum_{k}\Bigl{(}-\log\pi+\log\gamma^{(p)}_{\beta,k}-{|\beta^{(p)}_{k}|}^{2}\gamma^{(p)}_{\beta,k}$
$\,\quad\quad+(c_{0}-1)\log\gamma^{(p)}_{\beta,k}-d_{0}\gamma^{(p)}_{\beta,k}$
$\,\quad\quad+c_{0}\log d_{0}-\log(\Gamma(c_{0}))\Bigr{)}\,.$ (46)

### v2

@@ Abstract
Algorithms for mutual interference mitigation and object parameter estimation are a key enabler for automotive applications of frequency-modulated continuous wave (FMCW) radar. In this paper, we introduce a signal separation method to detect and estimate radar object parameters while jointly estimating and successively canceling the interference signal. The underlying signal model poses a challenge, since both the coherent radar echo and the non-coherent interference influenced by individual multipath propagation channels must be considered. Under certain assumptions, the model is described as a superposition of multipath channels weighted by parametric interference chirp envelopes. Inspired by sparse Bayesian learning (SBL), we employ an augmented probabilistic model that uses a hierarchical Gamma-Gaussian prior model for each multipath channel. Based on this, an iterative inference algorithm is derived using the variational expectation-maximization (EM) methodology. The algorithm is statistically evaluated in terms of object parameter estimation accuracy and robustness, indicating that it is fundamentally capable of achieving the Cramer-Rao lower bound (CRLB) with respect to the accuracy of object estimates and it closely follows the radar performance achieved when no interference is present.
@@ I-B Contributions and Notations
Although the existing literature is extensive, we identify the following open topics for which our work provides novel insights. Previously proposed algorithms are applied as preprocessing in a conventional range-Doppler processing chain. This approach is practical, but has inherent limitations. In general, the goal of radar sensing is to provide estimates of object parameters, which is equivalent to estimating the parameters of a multipath propagation channel containing static (such as buildings, traffics signs or trees) and moving objects (such as vehicles, bicycle or pedestrians). Note that parametric channel models typically represent multipath propagation as a linear superposition of weighted Dirac delta distributions - or spectral lines - with distinct supports in the underlying dispersion domain (range, angle of arrival, angle of departure, Doppler frequency, and combinations thereof). Therefore, estimating multipath parameters can be cast as a line spectral estimation (LSE) problem [56], for which SBL [57, 58] and related algorithms have been developed. An SBL-type method has been previously proposed as a preprocessing step to interpolate the received signal after interference zeroing [30], as referenced in Section I-A. Our proposed algorithm’s design and its aims are however distinct, as detailed in the remainder of this paper.
To achieve robust inference performance, it is necessary to consider the entire problem of parameter estimation under the influence of interfering signals. Previous methods are also often based on heuristics or require the prior setting of certain algorithm parameters. This makes it necessary to manually fine-tune the range-Doppler processing chain and makes it difficult to give performance guarantees. Our proposed algorithm is derived within a rigorous theoretical framework. Specific contributions are the following:
@@
  • We propose a novel inference algorithm inspired by SBL [58, 59, 60] for LSE that is able to infer the superposition of sparse line spectra. The algorithm is based on the variational EM approach. It jointly estimates the objects’ range-Doppler parameters with the parameters of the interference signal and multipath channel. This makes it explicitly robust to mutual interference. Note that the object parameters are estimated super-resolved in a grid-less manner [56].
@@
  • We include investigations on difficult-to-handle scenarios of model mismatch and poor signal separability. We point out the relative resilience of the algorithm to these challenges, and provide an initial proof of concept based on measurement data.
@@ II-A FMCW Radar under Interference
where we defined the parametrized rectangular function template $\bar{x}(t;T_{\mathrm{sw}})$ and the chirp $u(t;f_{0},k)$. $A$ is some complex transmit amplitude, $P$ is the number of chirps, $f_{0}$ is the chirp start frequency, $k$ is the chirp slope and $T_{\mathrm{p}}$ is the pulse duration which consists of the actual transmit chirp duration $T_{\mathrm{sw}}$ plus an additional idle time. [^2] Note that actual transmit signals are real-valued but a complex receiver architecture is assumed here. Thus, we immediately use a complex representation for mathematical simplicity. However, it is possible to model a real receiver architecture as well.
@@ III-D Algorithm Implementation
We have now derived structured mean-field variational EM update steps for inference of our model. As mentioned, these steps are interdependent and therefore to be applied iteratively. Explicit forms of the update equations are listed in Appendix B. The algorithm converges to an optimum of the ELBO corresponding to an approximate solution of our signal separation problem. However, it does not necessarily converge to the global optimum, nor is any rate of convergence guaranteed. In addition to the specific update schemes presented previously, the overall scheduling of steps as well as the initialization significantly influence the behavior of the algorithm.
@@
Received signal $\bm{r}$; thresholds $T_{\alpha}$, $T_{\beta}$
$\triangleright$ Bottom-up dictionary initialization:
$\hat{\bm{\Phi}}\leftarrow[\,]$, $\hat{\bm{\gamma}}_{\alpha}\leftarrow[\,]$, $\hat{\bm{\alpha}}\leftarrow[\,]$, $\hat{\bm{C}}_{\alpha}\leftarrow[\,]$
$\hat{\bm{\Psi}}\leftarrow[\,]$, $\hat{\bm{\gamma}}_{\beta}\leftarrow[\,]$, $\hat{\bm{\beta}}\leftarrow[\,]$, $\hat{\bm{C}}_{\beta}\leftarrow[\,]$
$\triangleright$ Noise precision initialization:
$\hat{\lambda}\leftarrow(2(NP-1))/{\|\bm{r}\|}^{2}$
@@
Algorithm 1 describes the implementation of the proposed iterative algorithm. For notational simplicity, we consistently omit iteration counters; within the sequence of update steps, the algorithm uses the most recent values of other estimates. The proposed algorithm structure envisions a bottom-up scheme, i.e., that the dictionaries $\hat{\bm{\Phi}}$ and analogously $\hat{\bm{\Psi}}$ are initialized as empty. The noise precision is initialized as $\hat{\lambda}=(2(NP-1))/{\|\bm{r}\|}^{2}$, i.e., half the received signal power is assumed to be from noise initially.
The main iteration consists of two separated subroutines for the estimations of the respective object and interference signals, given in Algorithms 2 and 3. They are similar in structure and connected by computations of $\hat{\bm{r}}_{\alpha}$, $\hat{\bm{r}}_{\beta}$, as well as $\hat{\lambda}$. I.e., the estimated residuals are used at every iteration as inputs to the respective subroutines. This scheme is obtained as the interference and object signals are additive and their components factorize according to (14) and (20). Furthermore, as interference is modeled independently over the ramps, separate steps over the index $p$ are applied. The estimation of multipath channels is based on the repeated addition and then update of line spectral components, in accordance with the bottom-up scheme, by applying (22) and Algorithm 4. The routine adds components it can “find” given the selected threshold, which are then refined or pruned using the now updated estimate of the weight covariance matrix.
@@ IV-A Fundamental Object Estimation Performance
Parameter Simulation I Simulation II Obj. Measurement Int. Measurement
@@ IV-C Statistical Performance Comparison
Figure 5: Statistical results of multi-object simulation scenario. (a)-(c) Sample mean over varying SNR at a fixed SIR of $-20\mathrm{dB}$. (d)-(f) Sample mean over varying SIR at a fixed SNR of $60\mathrm{dB}$. (g)-(i) Distribution of results at an SNR of $60\mathrm{dB}$ and SIR of $15\mathrm{dB}$ for each compared method, depicted in the form of box plots. Markers indicate outlier values, and the total number of outlier samples is additionally annotated.
@@
    mca - using the concept and model of [49]. Our implementation follows that of [73, Algorithm 30].
Figs. 5a-c depict the mean of the resulting metrics computed from 500 realizations over a varying SNR. Interference leads to a lower limit on the achievable estimation error and the number of misdetections, and causes false alarms. With preprocessing, the estimation error and number of misdetections improve, but the performance is still limited and improvements come at the cost of false alarms. This behavior can be explained by considering the effects of preprocessing on the form of the resulting signal. Zeroing merely exploits the time-limited nature of the interference, and hence leads to a signal with missing clusters of samples. MCA considers a signal separation problem, but does not employ a parametric model for the interference bursts and only applies single-ramp on-grid LSE for the object signal. Preprocessing therefore typically causes a distortion of the object signal that is a model error for the subsequent detector, primarily leading to false alarms. Our proposed method mitigates this behavior and reaches performance very close to the No Interference case. The estimation error and the number of misdetections decrease with increasing SNR, and the number of false alarms is constant according to the set detection threshold.
Figs. 5d-f show mean results at a high SNR, over SIR levels ranging from highly interfered to weakly interfered object signals. Ideally, perfect interference cancellation leads to values identical to the No Interference case at $\mathrm{SNR}=60\mathrm{dB}$. The proposed method approximately achieves this up until $\mathrm{SIR}\approx 0\mathrm{dB}$. Its performance then slightly decreases with decreasing interference power up until $\mathrm{SIR}\approx 15\mathrm{dB}$, after which it slowly recovers. This is the same effect as discussed in Section IV-A. Note that unlike for the simulation scenario of Fig. 3, in the scenario of Fig. 5 the condition of (27) is met roughly at $5\mathrm{dB}<\mathrm{SIR}<10\mathrm{dB}$. This strengthens our hypothesis on the significance of this condition. For the other evaluated methods, we see a bounded improvement in RMSE and misdetection rate at the cost of further false alarms. Naturally, perfect zeroing is independent of interference power while MCA performance is generally improved with increasing SIR.
Figs. 5g-i illustrate the distributions of the results at $\mathrm{SNR}=60\mathrm{dB}$ and $\mathrm{SIR}=-15\mathrm{dB}$ using box plots, in order to evaluate the robustness of the tested methods. In particular for the proposed method, this further gauges the risk of the algorithm converging to a local optimum of the signal separation problem, which could significantly decrease object detection performance. We indeed observe some outliers with higher RMSE values and an increased number of false alarms. However, the number of significant outliers, particularly when noting that the comparison is to be made to the ideal No Interference test case, is very small.
@@
IV-D Measurement Examples
Figure 7: Two examples of the proposed algorithm applied to real automotive radar measurement data. The first row is from the authors’ own combined object and interference measurements, while the second row is from a publicly available test dataset. White shaded rectangles indicate an object of interest that becomes undetectable due to interference. (a)/(e) Photo of the measurement scene. (b)-(c)/(f)-(g) Range-Doppler spectrum of the measured signal without/with interference and corresponding object detections. (d)/(h) Range-Doppler spectrum of the interference-cancelled signal and corresponding object detections as a result of the proposed algorithm.
@@ IV-D Measurement Examples
In order to illustrate the practical application of the proposed method, we apply it to radar measurement data from two different datasets. We select a single measurement each and showcase the results qualitatively; note that no ground truth of the channel parameters are available.
The first example was obtained as part of a measurement campaign in real traffic scenarios, details of which can be found in previous works [74, 75]. Measurement parameters are summarized in Table I (Obj. Measurement). These measurements do not inherently contain any interference. In prior work, simulated interference was added to the measured signals. In this work however, measured interference is added, as described in Appendix C, with parameters listed in Table I (Int. Measurement). The interference was obtained from a separate measurement, due to the difficulty of generating and measuring interference separately in the same inner city traffic environment. Nevertheless, the applied scheme enables a qualitative comparison of our mitigation result to an interference-free one, with both object and interference components obtained using actual radar sensors.
Fig. 7a shows a photo of the selected Obj. Measurement scene. Particularly, visible on the photo is the car in front and to the left of the ego-radar, as well as the facade of a row of buildings to the right. The interference-free spectrum in Fig. 7b hence contains a few close object components with a positive relative velocity, i.e., the car moving away from the sensor. The buildings lead to a large number of components distributed over distances, with negative relative velocities with respect to the measurement vehicle. Other detected components are reflections not easily associated with the camera image or false alarms due to noise. When strong interference is introduced, the range-Doppler spectrum becomes dominated by it, see Fig. 7c. Directly applying object detection onto this signal leads to no discernibly useful result. Finally, the spectrum of the measurement when subtracting the interference estimated by the proposed processing algorithm, and the estimated object components are shown in Fig. 7d. The resulting spectrum is visually very close to the interference-free one and the detected components correspond well to the significant parts of the described scene. We also qualitatively illustrate the fast-time signal of a single ramp in Fig. 8 as an example. The top graph shows a large interference burst not too dissimilar to, but distinct from the simulated ones of Figs. 2 and 4. The bottom plot provides a close-up view of the signal after subtraction of the interference estimate as well as shows the object signal estimate.
The second example measurement originates from [76, 77]. Measurement parameters and further details are found in the cited references. The measurement hardware and scenario are distinct from the previously presented measurement, in particular with respect to the interference. Here, the measurements inherently contain interference from two different sensors, so that no additional processing as described in Appendix C is necessary. The scenario is mostly static with significant reflections up to a distance of $20\mathrm{m}$, as well as a small drone flying around. Figs. 7e-h show the results for this example in an analogous manner to Figs. 7a-d. No exact ground truth is available, but the “clean” reference of Fig.7f is taken from the previous measurement frame, during which no interferer crossed the victim radar’s IF band. Similarly to the previous example, we can see that the mitigated spectrum as well as the detected components closely match the results for the interference-free measurement.
These very promising results are nevertheless only preliminary, and a larger scale analysis with better known hardware and measurement environment is a necessary part of future work.
@@ V Conclusion
Specific choices in the proposed algorithm design, initialization and scheduling are essential for its performance. This includes employing grid-less estimation for the object channel, yielding super-resolution accuracy, while using an on-grid algorithm for the interference channels. Parameter optimization and inference of the respective channel weights are derived from a joint update, which decreases the probability of converging to a local optimum of the signal separation problem.
@@ Appendix B Mean-field Variational EM Updates
We proceed to summarize the solution steps needed to obtain the update equations for our model. To clarify the mathematical derivations to follow, we write out the logarithms of factor distributions in the joint PDF (14), i.e. (15) to (18), explicitly so that
$\log p(\bm{r}|\bm{\alpha},$ $\bm{\beta},\lambda;\bm{\zeta},\bm{\theta})$
@@
$\,\quad-\lambda{(\bm{r}-\bm{\Phi\alpha}-\bm{U\Psi\beta})}^{\mathrm{H}}(\bm{r}-\bm{\Phi\alpha}-\bm{U\Psi\beta})$ (46)
$\log p{(\bm{\alpha}|\bm{\gamma}_{\alpha})}$ $=\sum_{l}\Bigl{(}-\log\pi+\log\gamma_{\alpha,l}-{|\alpha_{l}|}^{2}\gamma_{\alpha,l}\Bigr{)}$ (47)
$\log p{(\bm{\gamma}_{\alpha})}$ $=\sum_{l}\Bigl{(}(a_{0}-1)\log\gamma_{\alpha,l}-b_{0}\gamma_{\alpha,l}$
$\,\quad\quad+a_{0}\log b_{0}-\log(\Gamma(a_{0}))\Bigr{)}$ (48)
$\log p{(\bm{\beta}|\bm{\gamma}_{\beta})}$ $=\sum_{p}\sum_{k}\Bigl{(}-\log\pi+\log\gamma^{(p)}_{\beta,k}-{|\beta^{(p)}_{k}|}^{2}\gamma^{(p)}_{\beta,k}\Bigr{)}$ (49)
$p{(\bm{\gamma}_{\beta})}$ $=\sum_{p}\sum_{k}\Bigl{(}(c_{0}-1)\log\gamma^{(p)}_{\beta,k}-d_{0}\gamma^{(p)}_{\beta,k}$
$\,\quad\quad+c_{0}\log d_{0}-\log(\Gamma(c_{0}))\Bigr{)}$ (50)
$p(\lambda)$ $=(e_{0}-1)\log\lambda-f_{0}\lambda$
$\,\quad\quad+e_{0}\log f_{0}-\log(\Gamma(e_{0}))\,.$ (51)
We refer back to Section II for the introduction of this probabilistic model and its components.